OpenStudy (0487308):

Micah can fill a box with books in 15 minutes. While Micah fills the box, Sydney takes the books out puts them on a shelf. With Sydney taking books out of the box, it takes 20 minutes for Micah to fill the box. Which of the following can be used to determine the amount of time it takes for Sydney to empty the box if Micah is not adding books? Is the answer to this 1/15 - 1/x = 1/20? @kendricklamar2014

OpenStudy (mathmale):

Would you mind explaining your reasoning? With that info in mind, others may find it easier to give you appropriate feedback.

OpenStudy (0487308):

My reasoning is 1/15 because it is 1 box per 15 minutes, but then Sydney subtracts 1 box per x minutes, and you have 1 box per 20 minutes.

OpenStudy (0487308):

@mathmale

OpenStudy (mathmale):

I don't think that your "1/15 - 1/x = 1/20 " alone is sufficient. Instead, I'd write

OpenStudy (mathmale):

\[\frac{ 1.box.full }{ \frac{ 1.box }{ 15.\min }-\frac{ 1.box }{ x }}=20.\min\]

OpenStudy (mathmale):

Best to include the units of measurement always. Note that \[\frac{ 1.box.full }{ \frac{ 1.box }{ 15.\min }-\frac{ 1.box }{ x }}\] has units of minutes; "box" cancels out.

OpenStudy (0487308):

Okay, but none of those are answer choices. I only have 1/15 + 1/20 = 1/x 1/20 - 1/x = 1 / 15 1 / 15 - 1 / x = 1 / 20 1 / 15 - 1 / x = x / 20

OpenStudy (mathmale):

There are four equations here. It'd take only a few minutes to solve each one for x. That's what I'd recommend you do at the moment. Seems as though Sydney doesn't work very fast, since if he's emptying the box while M. fills it, M has to work only a little bit longer to fill the box. So, it seems to me that x would be larger than 15. This type of estimation can be valuable.

OpenStudy (mathmale):

the problem doesn't ask you to find x, true, but rather to choose the equation that will give you the correct value of x.

OpenStudy (0487308):

So it's either B or C?

OpenStudy (mathmale):

Sorry, 048, but I don't respond to "It's b" type questions. Need to see your work and to understand your reasoning.

OpenStudy (0487308):

Okay, my reasoning was above near the beginning of the comments for me thinking it's b. I don't see how it can be any of the other solutions.

OpenStudy (mathmale):

I've just tested one of those four choices and came up with x=60 minutes (1 hour). Again, we're not req'd to find x itself, but rather to choose the appropriate equation. challenge question for you: How could you test your choice of one of those four questions?

OpenStudy (0487308):

I do not need to solve for x in the situation, but I can see which ones make sense. I already submitted my exam and got a 100% on it.

OpenStudy (mathmale):

So glad you have. Congrats. Over and out.

OpenStudy (0487308):

Okay.

OpenStudy (mathmale):

Bye for now.